OpenStudy (anonymous):

Multiply and Simplify. √2( √8 + 4√3) im starting to understand these but the problem looks different so its hard to solve it

OpenStudy (anonymous):

\[\sqrt{2}(\sqrt{8}+4\sqrt{3})=\sqrt{2\times 8}+4\sqrt{3\times 2}\] \[=\sqrt{16}+4\sqrt{6}=4+4\sqrt{6}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw